OpenStudy (lgbasallote):

what is the yp in \[(D^4 + 2D^3 + 5D^2)y = x^2 - e^{-2x}\] i got \[y_p = \frac{e^{-2x}}{20}\] or something like that

OpenStudy (anonymous):

solve for \(y_c\) \[ D^4+2D^3+5D^2=D^2(D^2+2D+5)=0 \] it gives \(D=0,0, -1\pm2i \) so \(y_c=c_1+c_2x+e^{-x}(c_3\cos 2x+c_4 \sin 2x ) \) now for \(y_p\) see \(R(x)=x^2-e^{-2x}\) for \(x^2\) u should apply \(ax^2+bx+c\) but \(bx+c\) is similiar to \(c_1x+c_2\) in \(y_c\) so u must multiply \(ax^2+bx+c\) by \(x^2\) for Removing similarity...for \(e^{-2x}\) u must apply \(de^{-2x}\) so \(y_p=x^2(ax^2+bx+c)+de^{-2x}\) plug it in the original equation and find a,b,c and d final answer: \[y_p=x^2(\frac{25}{2}x^2-20x-6)-\frac{75}{2} e^{-2x} \]

OpenStudy (lgbasallote):

im not familiar witht he method you used o.O whatever it is...

OpenStudy (lgbasallote):

i just know undetermined coefficients and variation of paramaters...and that doesnt look like either

OpenStudy (anonymous):

variation of paramaters will work here...

OpenStudy (lgbasallote):

i think undetermined coefficients will too...can you show me the process using that?

OpenStudy (anonymous):

sorry igba almost forgot them...

OpenStudy (anonymous):

and wait actually what i did is undetermined coefficients i think...

OpenStudy (lgbasallote):

really? they dont look familiar to me o.O

OpenStudy (anonymous):

@lgbasallote what mukushla did is method of undetermined coefficients. just take the respective derivative \[y_{p}=x^2(a^2+bx+c)+de^{-2x}\] and put in the original equation then compare the coefficients .

OpenStudy (lgbasallote):

what the heck is that ax^2 + bx + c? o.O they didnt teach that to us in undetermined coefficients

OpenStudy (anonymous):

what they taught you how to take guess for polynomial?

OpenStudy (lgbasallote):

kind of....well they just taught us the two basic forms \[c_1 e^{mx}\] and then \[e^{ax} (c_1 \cos bx + c_2 \sin bx)\] and then the yp is whichever looks like one of those

OpenStudy (anonymous):

some guess for particular functions are given below.Taken from Paul online DE notes

OpenStudy (lgbasallote):

so i took a guess and i got \[Ax^2 - Be^{-2x}\]

OpenStudy (anonymous):

i think the following example will be good to understand .taken from Paul online notes .

OpenStudy (lgbasallote):

\[Dy_p = 2Ax + 2Be^{-2x}\] \[D^2y_p = 2A - 4Be^{-2x}\] \[D^3 y_p = 2 + 8Be^{-2x}\] \[D^4 y_p = -16e^{-2x}\]

OpenStudy (lgbasallote):

\[-16e^{-2x} +2(2 + 8De^{-2x}) + 5(2A - 4Be^{-2x}) = x^2 - e^{-2x}\]

OpenStudy (lgbasallote):

\[4 + 10A - 20 Be^{-2x} = x^2 - e^{-2x}\]

OpenStudy (lgbasallote):

\[A = 0\] \[B = \frac{1}{20}\]

OpenStudy (lgbasallote):

that's my solution

OpenStudy (anonymous):

whenever you have something like polynomial we generally write the full form of that degree polynomial for example if you have x^2 term in the right side then in guess for yp you must write \[y_{P}=ax^2+bx+c\] (general form of second degree polynomial) if you have x^3 in the write side then \[y_{p}=ax^3+bx^2+cx+d\] (general form of third degree polynomial) if you have x^4 in the write side then \[y_{p}=ax^4+bx^3+cx^2+dx+e\] (general form of 4th degree polynomial) i hope now you can guess it correctly for polynomials !

OpenStudy (lgbasallote):

...still not making sense....

OpenStudy (anonymous):

it should make sense now because you have x^2 terms in the right hand side and you must use yp guess following \[y_{p}=ax^2+bx+c+de^{-2x}\] but to remove duplication (terms are also in the complementary solution) multiply with x^2 as mukushla did you are using guess \[Ax^2-Be^{-2x}\] which is incorrect

OpenStudy (lgbasallote):

so bottomline here is im incorrect and i might have failed the "easiest" quiz in differential equations??!!! o.O

OpenStudy (anonymous):

i am sorry to say yes:(

OpenStudy (lgbasallote):

<--going into avatar state now

OpenStudy (anonymous):

do not worry you wouldn't repeat this in your Examinations after this :) what if the polynomial appears in the final exam :! m sure you will be able to do it better than any one :)

OpenStudy (lgbasallote):

too late to stop the avatar state >:(